# Generative Medical Event Models Improve with Scale

### v1

@@ 2.3.1 Disease-Specific Outcome Predictions
We measured CoMET’s performance on relevant disease-specific outcome prediction tasks, indexed to the time of care decisions. In particular, for type 2 diabetes mellitus (T2DM), hyperlipidemia (HLD), and hypertension (HTN), we examined CoMET’s ability to predict patients’ risk of relevant outcomes at the time of a change in medication management. Adverse outcomes, such as three-year stroke risk, were labeled as binary targets indicating whether the event occurred in the given time frame. Relevant lab results, such as hemoglobin A1c (HgbA1c) or total cholesterol, were labeled as binary targets at certain thresholds (e.g., HgbA1c <7<\!7).
@@
Figure 5 illustrates CoMET’s predictive performance on relevant outcomes for patients initiating a new medication therapy for active T2DM. These include outcomes such as one-year and three-year risk for atherosclerotic cardiovascular disease (ASCVD), chronic kidney disease (CKD) progression from stage 2 through stage 4+, diabetic neuropathy, and diabetic retinopathy, as well as two-to-four-month HgbA1c lab results (the time frame when this lab is recommended to be reassessed after medication changes [20]). CoMET models improved consistently in these discriminative prediction tasks, with CoMET-L outperforming task-specific supervised models on most of these tasks. Scatter plots for predicted lab values can be found in Figure 24.
@@
Figure 5: T2DM-specific outcome predictions. Percent increase of area under the curve of the receiver operating characteristic curve (AUCROC) from the best-performing task-specific supervised model for each of the three CoMET models on the T2DM-specific outcome prediction tasks.
@@
Likewise, Figure 6 shows results for CoMET predictions on tasks relevant to patients receiving treatment for hyperlipidemia, including one- and three-year risk of ASCVD, heart attacks, strokes, and chronic heart failure (e.g., only the chronic diagnosis codes related to chronic heart failure). CoMET-L achieves an AUCROC of 0.93 for predicting chronic heart failure diagnosis events within a year of changing hyperlipidemia medical management. While CoMET-L performance did not exceed the task-specific supervised models’ performance, CoMET models showed measurable improvements as they scaled up, and absolute AUCROC values were robust. Notably, CoMET did not outperform the task-specific models on the hyperlipidemia outcomes; the factors underlying this underperformance remain unclear and are left to future work.
@@
Figure 6: Hyperlipidemia-specific outcome predictions. Percent increase of AUCROC from the best-performing task-specific supervised model for each of the three CoMET models on the hyperlipidemia-specific outcome prediction tasks. CoMET models consistently scored better with scale, yet CoMET-L scored lower than the supervised models on each diagnosis task.
@@
Figure 7: Hypertension-specific outcome predictions. Percent increase of AUCROC from the best-performing task-specific supervised model for each of the three CoMET models on the hypertension-specific outcome prediction tasks. CoMET-L scores matched or exceeded the supervised models on 6 out of 10 tasks.
@@
For hypertension-related outcomes, we assessed CoMET performances to make predictions about one- and three-year risk of ASCVD events, heart attack, stroke, and CKD progression from stage 2 to stage 3 and from stage 3 to stage 4+. As with the two above cases, CoMET models improved as they increased in scale, with CoMET-L achieving higher AUCROC scores than the task-specific models on half of these tasks. A full list of evaluation scores across disease-specific outcome tasks can be found in Table 8.
@@ 2.3.2 Acute-on-Chronic Incidence Prediction
We next evaluated CoMET’s ability to predict the two-year risk of acute-on-chronic clinical events, such as asthma exacerbation or sickle cell crisis. Patients are included in each acute-on-chronic cohort if they demonstrate a medical history of the relevant chronic disease (e.g., sickle cell disease for sickle cell crisis).
We formulated each acute-on-chronic evaluation as a binary classification task: for congestive heart failure (CHF) exacerbations for patients with chronic CHF, asthma attacks for patients with asthma, sickle-cell crises for patients with sickle cell disease, alcohol withdrawal syndrome for patients with alcohol use disorder, and chronic obstructive pulmonary disease (COPD) exacerbations for patients with COPD. Detailed phenotype definitions, prediction date selection, and distinctions between chronic and acute event coding are provided in Section 5.6.5. Approximately 5,000 unique patients are included for each group, with optional upsampling to ensure at least 500 patients experienced the acute event within two years of their prediction date. Dataset characteristics for each task can be found in Table 19.
Figure 8: Acute-On-Chronic Tasks. Percent increase of AUCROC for each of the three CoMET models compared with the best-performing task-specific supervised model on the acute-on-chronic outcome prediction tasks. CoMET-M and CoMET-L scored higher than the baselines on 4 out of 5 tasks.
CoMET-L achieved a higher AUCROC than the task-specific models across all but one of these acute-on-chronic tasks, as shown in Figure 8, and in all tasks measured by PR-AUC. All tasks showed consistently increasing AUCROC and PR-AUC with larger CoMET model size. Tabulated results can be found in Table 9.
@@ 2.3.3 Incident Disease Risk Prediction
Finally, we tested CoMET’s performance on predicting the first instance of a variety of disease states within a two-year period in the general population. CoMET generated 60 simulated timelines for approximately 5,000 patients, with each generation spanning two years of tokens. To avoid cohort biases and priors, broad inclusion and minimal exclusion criteria were used on these task cohorts (see Section 5.6.4 for details). Notably, this leads to class imbalance where all targets have a positive prevalence <1.5%. This is quite different from the two preceding disease outcomes tasks, where the inclusion criteria for each task meant that those patients were naturally at increased risk for the prediction targets.
@@
Figure 9 compares CoMET to supervised models on incident disease tasks. AUCROC scores are reported in the main figure for consistency with the other disease risk tasks. CoMET demonstrated higher AUCROC scores than the supervised models on one out of six tasks and shows improvement across most tasks with model scale. However, in tasks like incident disease prediction with highly imbalanced class labels, PR-AUC is more commonly used to judge performance [21, 22]. CoMET-L achieved higher PR-AUC scores than the supervised models on all six tasks. Results for CoMET and supervised models (including PR-AUC) are reported in Table 10. One possible reason that CoMET-L did not outperform most task-specific models in AUCROC is because these tasks had extremely low prevalence and more generations per patient may be necessary. Preliminary evidence for this can be seen in Section 2.7.2.
@@
Figure 9: Incident Disease Risk Predictive Performance. Percent increase of AUCROC from the best-performing task-specific supervised model for each of the three CoMET models on the six incident-disease prediction tasks. CoMET-L scores higher than the baselines on one out of six tasks.
2.4 CoMET models generate early, quantitative differential diagnoses
@@
We next assessed whether CoMET can provide ranked, quantitative differential diagnoses for individual patients. We selected two clusters of diagnoses, hepatopancreatobiliary (HPB) diseases and rheumatic diseases, because they span a variety of clinical presentations, can take multiple encounters to diagnose definitively, and have challenges with delayed diagnosis or misdiagnosis [23, 24, 25, 26, 27, 28, 29, 30]. Diseases within each cluster often present clinically with overlapping signs, symptoms, and laboratory findings, and as a result they often appear together on differential diagnoses. We selected nine diagnoses for both HPB and rheumatic diseases, and for each we selected a cohort of patients who received that diagnosis (see Section 5.6.6 for diagnoses and their code-based definitions, and Table 20 for sample sizes). For each patient, we chose several points in their history in the one-year span before the first occurrence of their target diagnosis and had CoMET predict their risk for all the cluster’s diagnoses. In this way, CoMET produced a ranked, quantitative differential diagnosis at multiple time points for each patient and flagged patients at risk of receiving these diagnoses. We did not prompt CoMET with any information beyond the start of the encounter at which the patient first received their target diagnosis, so CoMET never sees any diagnostic workup or documentation from this encounter.
@@
Each plot in the top row of Figure 10 and Figure 11 shows how many patients were flagged by CoMET-L as having at least 10% risk for the target diagnosis (bold) and each of the off-target diagnoses within the cluster (lighter, thinner lines). Each plot therefore represents the average differential diagnosis for the cohort over time. For the majority of diagnoses, CoMET-L correctly flagged more than 50% of patients for their target diagnosis by their final prediction time. For most diagnoses, CoMET-L also flagged more than 25% of patients weeks ahead of time.
@@
Figure 10: Hepatopancreatobiliary differential diagnosis. CoMET-L was evaluated on predicting 1.5-year risk of receiving one of the indicated HPB diagnoses. Each column represents a cohort of patients who were ultimately diagnosed with the indicated diagnosis. Each line in the first row represents the percentage of that cohort that was flagged by CoMET-L as having at least 10% risk of a diagnosis. The correct diagnosis is shown in bold, the off-target diagnoses are faint; each line color represents the same diagnosis across the row. The second row shows the AUCROC over time for all three CoMET models and the task-specific supervised model for predicting the indicated diagnosis.
@@
Figure 11: Rheumatic differential diagnosis. CoMET-L was evaluated on predicting 1.5-year risk of receiving one of the indicated rheumatic diagnoses. Each column represents a cohort of patients who were ultimately diagnosed with the indicated diagnosis. Each line in the first row represents the percentage of that cohort that was flagged by CoMET-L as having at least 10% risk of a diagnosis. The correct diagnosis is shown in bold, the off-target diagnoses are faint; each line color represents the same diagnosis across the row. The second row shows the AUCROC over time for all three CoMET models and the task-specific supervised model for predicting the indicated diagnosis.
@@
For comparison, we trained task-specific supervised models to predict each of these diagnoses in a wider, more general pool of patients (see Section 5.4 for details on training). These task-specific models were generally not able to differentiate well among patients within these narrower HPB or rheumatic diagnosis clusters, with AUCROC metrics not increasing appreciably over time (see the bottom row of both Figure 10 and Figure 11), despite achieving modest sensitivity. In contrast, AUCROC metrics from CoMET generally increase over time and as model size increases. This suggests that CoMET models can effectively distinguish patients with similar presentations but different eventual diagnoses, and they can produce realistic differential diagnoses that become more sensitive and specific as the patient’s clinical presentation and diagnostic workup evolve. AUCROC at the final prediction time is reported for all diagnoses and models in Table 11. Of note, these analyses do not address whether early diagnosis flagging reflects the diagnostic workup being pursued by the patient’s medical providers versus the ability to preemptively flag diagnoses before significant clinical suspicion.
@@
Figure 10 shows that, among all nine HPB diseases, the correct diagnosis was the one most commonly flagged by CoMET-L, with the gap between the first- and second-ranked diagnoses generally increasing with time. For acute pancreatitis, AUCROC is hardly better than chance until the target diagnosis date. For diseases with generally more insidious onset (e.g., cancer, chronic viral hepatitis, and alcoholic liver disease), CoMET-L was able to flag many patients much earlier (Figure 10).
@@
For rheumatic disorders (Figure 11), most patients were flagged at the 10% risk levels for osteoarthritis at some point, regardless of their eventual diagnosis. This is likely both because osteoarthritis is a common diagnosis for patients with undifferentiated joint pain and because of early inaccurate or imprecise diagnosis and documentation. Among the remaining eight target diagnoses, the correct diagnosis was ranked second for five cohorts, third for two, and fourth for one. AUCROC scores for CoMET models ranged from 0.66-0.79 across rheumatic diagnoses at the final prediction time (Table 11).
@@
2.5 CoMET models forecast patients’ interactions with the health system
@@
Reliably forecasting health system interactions enables clinicians and health systems to plan for the needs of their patients. We assessed CoMET’s generated patient timelines for their ability to make accurate predictions about a patient’s interactions with the healthcare system.
@@
Having earlier assessed the probability calibration of predicting the number of encounters over a year in Figure 3, we next asked how close these predictions were to the ground truth number of encounters. For inpatient, outpatient, and emergency encounters, all three CoMET models demonstrated lower mean absolute error (MAE) than supervised task-specific regression models for predicting future encounter counts. Results are shown in Figure 12.

### v3

@@ 2.3.1 Disease-Specific Outcome Predictions
We measured Curiosity’s performance on relevant disease-specific outcome prediction tasks, indexed to the time of care decisions. In particular, for type 2 diabetes mellitus (T2DM), hyperlipidemia (HLD), and hypertension (HTN), we examined Curiosity’s ability to predict patients’ risk of relevant outcomes at the time of a change in medication management. Adverse outcomes, such as three-year stroke risk, were labeled as binary targets indicating whether the event occurred in the given time frame. Relevant lab results, such as hemoglobin A1c (HgbA1c) or total cholesterol, were labeled as binary targets at certain thresholds (e.g., HgbA1c <7<\!7).
@@
Figure 5 illustrates Curiosity’s predictive performance on relevant outcomes for patients initiating a new medication therapy for active T2DM. These include outcomes such as one-year and three-year risk for atherosclerotic cardiovascular disease (ASCVD), chronic kidney disease (CKD) progression from stage 2 through stage 4+, diabetic neuropathy, and diabetic retinopathy, as well as two-to-four-month HgbA1c lab results (the time frame when this lab is recommended to be reassessed after medication changes [20]). Curiosity models improved consistently in these discriminative prediction tasks, with Curiosity-L outperforming task-specific supervised models on most of these tasks. Scatter plots for predicted lab values can be found in Figure 24.
@@
Figure 5: T2DM-specific outcome predictions. Percent increase of area under the curve of the receiver operating characteristic curve (AUCROC) from the best-performing task-specific supervised model for each of the three Curiosity models on the T2DM-specific outcome prediction tasks.
@@
Likewise, Figure 6 shows results for Curiosity predictions on tasks relevant to patients receiving treatment for hyperlipidemia, including one- and three-year risk of ASCVD, heart attacks, strokes, and chronic heart failure (e.g., only the chronic diagnosis codes related to chronic heart failure). Curiosity-L achieves an AUCROC of 0.93 for predicting chronic heart failure diagnosis events within a year of changing hyperlipidemia medical management. While Curiosity-L performance did not exceed the task-specific supervised models’ performance, Curiosity models showed measurable improvements as they scaled up, and absolute AUCROC values were robust. Notably, Curiosity did not outperform the task-specific models on the hyperlipidemia outcomes; the factors underlying this underperformance remain unclear and are left to future work.
@@
Figure 6: Hyperlipidemia-specific outcome predictions. Percent increase of AUCROC from the best-performing task-specific supervised model for each of the three Curiosity models on the hyperlipidemia-specific outcome prediction tasks. Curiosity models consistently scored better with scale, yet Curiosity-L scored lower than the supervised models on each diagnosis task.
@@
Figure 7: Hypertension-specific outcome predictions. Percent increase of AUCROC from the best-performing task-specific supervised model for each of the three Curiosity models on the hypertension-specific outcome prediction tasks. Curiosity-L scores matched or exceeded the supervised models on 6 out of 10 tasks.
@@
For hypertension-related outcomes, we assessed Curiosity’s performances to make predictions about one- and three-year risk of ASCVD events, heart attack, stroke, and CKD progression from stage 2 to stage 3 and from stage 3 to stage 4+. As with the two above cases, Curiosity models improved as they increased in scale, with Curiosity-L achieving higher AUCROC scores than the task-specific models on half of these tasks. A full list of evaluation scores across disease-specific outcome tasks can be found in Table 8.
@@ 2.3.2 Acute-on-Chronic Incidence Prediction
We next evaluated Curiosity’s ability to predict the two-year risk of acute-on-chronic clinical events, such as asthma exacerbation or sickle cell crisis. Patients are included in each acute-on-chronic cohort if they demonstrate a medical history of the relevant chronic disease (e.g., sickle cell disease for sickle cell crisis).
We formulated each acute-on-chronic evaluation as a binary classification task: for congestive heart failure (CHF) exacerbations for patients with chronic CHF, asthma attacks for patients with asthma, sickle cell crises for patients with sickle cell disease, alcohol withdrawal syndrome for patients with alcohol use disorder, and chronic obstructive pulmonary disease (COPD) exacerbations for patients with COPD. Detailed phenotype definitions, prediction date selection, and distinctions between chronic and acute event coding are provided in Section 5.6.5. Approximately 5,000 unique patients are included for each group, with optional upsampling to ensure at least 500 patients experienced the acute event within two years of their prediction date. Dataset characteristics for each task can be found in Table 19.
Figure 8: Acute-On-Chronic Tasks. Percent increase of AUCROC for each of the three Curiosity models compared with the best-performing task-specific supervised model on the acute-on-chronic outcome prediction tasks. Curiosity-M and Curiosity-L scored higher than the baselines on 4 out of 5 tasks.
Curiosity-L achieved a higher AUCROC than the task-specific models across all but one of these acute-on-chronic tasks, as shown in Figure 8, and in all tasks measured by PR-AUC. All tasks showed consistently increasing AUCROC and PR-AUC with larger Curiosity model size. Tabulated results can be found in Table 9.
@@ 2.3.3 Incident Disease Risk Prediction
Finally, we tested Curiosity’s performance on predicting the first instance of a variety of disease states within a two-year period in the general population. Curiosity generated 60 simulated timelines for approximately 5,000 patients, with each generation spanning two years of tokens. To avoid cohort biases and priors, broad inclusion and minimal exclusion criteria were used on these task cohorts (see Section 5.6.4 for details). Notably, this leads to class imbalance where all targets have a positive prevalence <1.5%. This is quite different from the two preceding disease outcomes tasks, where the inclusion criteria for each task meant that those patients were naturally at increased risk for the prediction targets.
@@
Figure 9 compares Curiosity to supervised models on incident disease tasks. AUCROC scores are reported in the main figure for consistency with the other disease risk tasks. Curiosity demonstrated higher AUCROC scores than the supervised models on one out of six tasks and shows improvement across most tasks with model scale. However, in tasks like incident disease prediction with highly imbalanced class labels, PR-AUC is more commonly used to judge performance [21, 22]. Curiosity-L achieved higher PR-AUC scores than the supervised models on all six tasks. Results for Curiosity and supervised models (including PR-AUC) are reported in Table 10. One possible reason that Curiosity-L did not outperform most task-specific models in AUCROC is because these tasks had extremely low prevalence and more generations per patient may be necessary. Preliminary evidence for this can be seen in Section 2.7.2.
@@
Figure 9: Incident Disease Risk Predictive Performance. Percent increase of AUCROC from the best-performing task-specific supervised model for each of the three Curiosity models on the six incident-disease prediction tasks. Curiosity-L scores higher than the baselines on one out of six tasks.
2.4 Curiosity models generate early, quantitative differential diagnoses
@@
We next assessed whether Curiosity can provide ranked, quantitative differential diagnoses for individual patients. We selected two clusters of diagnoses, hepatopancreatobiliary (HPB) diseases and rheumatic diseases, because they span a variety of clinical presentations, can take multiple encounters to diagnose definitively, and have challenges with delayed diagnosis or misdiagnosis [23, 24, 25, 26, 27, 28, 29, 30]. Diseases within each cluster often present clinically with overlapping signs, symptoms, and laboratory findings, and as a result they often appear together on differential diagnoses. We selected nine diagnoses for both HPB and rheumatic diseases, and for each we selected a cohort of patients who received that diagnosis (see Section 5.6.6 for diagnoses and their code-based definitions, and Table 20 for sample sizes). For each patient, we chose several points in their history in the one-year span before the first occurrence of their target diagnosis and had Curiosity predict their risk for all the cluster’s diagnoses. In this way, Curiosity produced a ranked, quantitative differential diagnosis at multiple time points for each patient and flagged patients at risk of receiving these diagnoses. We did not prompt Curiosity with any information beyond the start of the encounter at which the patient first received their target diagnosis, so Curiosity never sees any diagnostic workup or documentation from this encounter.
@@
Each plot in the top row of Figure 10 and Figure 11 shows how many patients were flagged by Curiosity-L as having at least 10% risk for the target diagnosis (bold) and each of the off-target diagnoses within the cluster (lighter, thinner lines). Each plot therefore represents the average differential diagnosis for the cohort over time. For the majority of diagnoses, Curiosity-L correctly flagged more than 50% of patients for their target diagnosis by their final prediction time. For most diagnoses, Curiosity-L also flagged more than 25% of patients weeks ahead of time.
@@
Figure 10: Hepatopancreatobiliary differential diagnosis. Curiosity-L was evaluated on predicting 1.5-year risk of receiving one of the indicated HPB diagnoses. Each column represents a cohort of patients who were ultimately diagnosed with the indicated diagnosis. Each line in the first row represents the percentage of that cohort that was flagged by Curiosity-L as having at least 10% risk of a diagnosis. The correct diagnosis is shown in bold, the off-target diagnoses are faint; each line color represents the same diagnosis across the row. The second row shows the AUCROC over time for all three Curiosity models and the task-specific supervised model for predicting the indicated diagnosis.
@@
Figure 11: Rheumatic differential diagnosis. Curiosity-L was evaluated on predicting 1.5-year risk of receiving one of the indicated rheumatic diagnoses. Each column represents a cohort of patients who were ultimately diagnosed with the indicated diagnosis. Each line in the first row represents the percentage of that cohort that was flagged by Curiosity-L as having at least 10% risk of a diagnosis. The correct diagnosis is shown in bold, the off-target diagnoses are faint; each line color represents the same diagnosis across the row. The second row shows the AUCROC over time for all three Curiosity models and the task-specific supervised model for predicting the indicated diagnosis.
@@
For comparison, we trained task-specific supervised models to predict each of these diagnoses in a wider, more general pool of patients (see Section 5.4 for details on training). These task-specific models were generally not able to differentiate well among patients within these narrower HPB or rheumatic diagnosis clusters, with AUCROC metrics not increasing appreciably over time (see the bottom row of both Figure 10 and Figure 11), despite achieving modest sensitivity. In contrast, AUCROC metrics from Curiosity generally increase over time and as model size increases. This suggests that Curiosity models can effectively distinguish patients with similar presentations but different eventual diagnoses, and they can produce realistic differential diagnoses that become more sensitive and specific as the patient’s clinical presentation and diagnostic workup evolve. AUCROC at the final prediction time is reported for all diagnoses and models in Table 11. Of note, these analyses do not address whether early diagnosis flagging reflects the diagnostic workup being pursued by the patient’s medical providers versus the ability to preemptively flag diagnoses before significant clinical suspicion.
@@
Figure 10 shows that, among all nine HPB diseases, the correct diagnosis was the one most commonly flagged by Curiosity-L, with the gap between the first- and second-ranked diagnoses generally increasing with time. For acute pancreatitis, AUCROC is hardly better than chance until the target diagnosis date. For diseases with generally more insidious onset (e.g., cancer, chronic viral hepatitis, and alcoholic liver disease), Curiosity-L was able to flag many patients much earlier (Figure 10).
@@
For rheumatic disorders (Figure 11), most patients were flagged at the 10% risk levels for osteoarthritis at some point, regardless of their eventual diagnosis. This is likely both because osteoarthritis is a common diagnosis for patients with undifferentiated joint pain and because of early inaccurate or imprecise diagnosis and documentation. Among the remaining eight target diagnoses, the correct diagnosis was ranked second for five cohorts, third for two, and fourth for one. AUCROC scores for Curiosity models ranged from 0.66-0.79 across rheumatic diagnoses at the final prediction time (Table 11).
@@
2.5 Curiosity models forecast patients’ interactions with the health system
@@
Reliably forecasting health system interactions enables clinicians and health systems to plan for the needs of their patients. We assessed Curiosity’s generated patient timelines for their ability to make accurate predictions about a patient’s interactions with the healthcare system.
@@
Having earlier assessed the probability calibration of predicting the number of encounters over a year in Figure 3, we next asked how close these predictions were to the ground truth number of encounters. For inpatient, outpatient, and emergency encounters, all three Curiosity models demonstrated lower mean absolute error (MAE) than supervised task-specific regression models for predicting future encounter counts. Results are shown in Figure 12.